What is the period for y = tan (3 pi / 4 theta) and give where the two asymptotes occur for the function

Answer by stanbon(75887)

What is the period for y = tan (3 pi / 4 theta) and give where the two asymptotes occur for the function
-----
Period = pi/(3pi/4) = 4/3 radians
----
Asymptotes where (3pi/4)theta = pi/2 or (3pi/4)theta = (3pi/2)
---
Asymptotes at theta = (4/3)(1/2) = 2/3 radians or theta = 2 radians
==============================
Cheers,
Stan H.